OpenStudy (anonymous):

Write the polynomial in factored form. 4x3 + 8x2 – 96x

OpenStudy (anonymous):

I need someone to explain to me how to do this.

OpenStudy (e.mccormick):

\(4x^3 + 8x^2 – 96x\) OK, you want things to add up, but one od the simple thing to look at first is if you can factor out a constant.

OpenStudy (anonymous):

you can't, right?

OpenStudy (e.mccormick):

Well, they are eve so you have to be able to factor out a 2, and possibly a 4.

OpenStudy (e.mccormick):

What do you get if you try to factor out a 4?

OpenStudy (e.mccormick):

And an x, since there is at least one x in each term.

OpenStudy (anonymous):

2 and 4 and 48? is that what you mean?

OpenStudy (anonymous):

4x3+8x2-96x=4x(x2+2x-24) =4x(x2+6x-4x-24) =4x{x(x+6)-4(x+6)} =4x(x+6)(x-4)

OpenStudy (anonymous):

ok I can kinda see where that came from

OpenStudy (e.mccormick):

Well, that would be if you factored out a 2. But if you factor out the 4, you get 1, 2, and -24. But surjit has just posted all the steps....

OpenStudy (anonymous):

thanks

OpenStudy (e.mccormick):

Have any questions on how that was done?

OpenStudy (anonymous):

not really sure were 24 came from

OpenStudy (e.mccormick):

96/4

OpenStudy (e.mccormick):

25*4=100. and 96 is 4 less than 100, so 24*4=96.

OpenStudy (anonymous):

and 8/4?

OpenStudy (e.mccormick):

Is the 2.

OpenStudy (anonymous):

and 4/4 is the x^2 i think i get it

OpenStudy (e.mccormick):

Yah, that is the way to factor out from all of them. Find something that divides out of all the terms. Then you separate it with () because it would multiply back to all the terms.

OpenStudy (anonymous):

gotchya thanks

OpenStudy (e.mccormick):

np. Have fun!